### v1

# 2022-deformable-elastic-objects-UAVs

Abstract
Aerial cooperative robotic manipulation of cable-suspended objects has been largely studied as it allows handling large and heavy objects, and cables offer multiple advantages, such as their low weight and cost efficiency. Multirotors have been typically considered, which, however, can be unsuitable for long-lasting manipulation tasks due to their scarce endurance. Hence, this work investigates whether non-stop flights are possible for maintaining constant the pose of cable-suspended objects. First, we show that one or two flying carriers alone cannot perform non-stop flights while maintaining a constant pose of the suspended object. Instead, we demonstrate that three flying carriers can achieve this task provided that the orientation of the load at the equilibrium is such that the components of the cable forces that balance the external force (typically gravity) do not belong to the plane of the cable anchoring points on the load. Numerical tests are presented in support of the analytical results.
I Introduction
Aerial robotic object manipulation has been largely studied also thanks to its interesting applications, e.g., transportation and assembly, to name a few [1]. Among the several manipulator designs that have been proposed in the literature [2], lightweight manipulation tools such as simple low-cost cables have attracted a lot of attention [3], especially due to the flying vehicle payload limitations.
Uncrewed helicopters have been used in [4] for transportation of suspended objects, while multirotor UAVs (Uncrewed Aerial Vehicles) have been typically considered [5, 6, 7, 8, 9, 10]. In the literature, we find examples of a slung load suspended below a single multi-rotor [11, 12] or a small helicopter [13]. Two multirotors have been widely exploited, especially for the manipulation of bar-shaped objects [14, 15, 16]. Three is the minimum number of robots to control the full pose of a cable-suspended rigid body object [17]; cooperative manipulation of cable-suspended load below three multirotors have been studied, e.g., in [18, 5, 7, 19].
Compared to multirotor UAVs, fixed-wing UAVs have a longer flight endurance, as it has been shown, e.g., in [20]. However, fixed-wing UAVs cannot stop in mid-air, thus making seemingly impossible to keep the load in a static equilibrium while using such platforms as flying carriers. Furthermore these platforms are unable of Vertical Take Off and Landing (VTOL). With an intermediate flight endurance, convertible UAVs such as tail sitters have been proposed [20]. These are VTOL platforms that convert themselves to proceed flying horizontally as a fixed-wing platform. However, also these designs are obliged to follow non-stop trajectories to ensure an efficient flights.
Therefore, while non-stop UAVs may provide better endurance for cooperative cable suspended transportation tasks than multirotor UAVs, to the best of the authors’ knowledge, it has not been systematically studied yet whether non-stop flights are compatible with keeping the load in a static equilibrium pose, i.e., whether it is possible to have at the same time the pose of the load being static while the flying carriers keep loitering above the object.
The theoretical discovery of such compatibility would pave the way to a completely new type of cable-suspended aerial systems which would enable a much more energy-efficient execution of aerial manipulation tasks in which long-distance transport is combined with fine manipulation of the object’s pose at the destination or in the presence of difficult intermediate passages, e.g., in cluttered environments. An abstract representation of the future scenario that would be possibly enabled by this study is shown in Figure 1.
Figure 1: Abstract representation of the concept whose theoretical feasibility is studied in this work. A team of non-stop flying carriers regulate the pose of a suspended object to a static forced equilibrium in an ideal application scenario.
In this work, we take up the challenge posed by this intriguing problem and we study analytically whether there exist non-stop trajectories for the flying carriers that are compatible with the constraint of keeping the load at a constant static pose. We focus on the case in which the load must maintain a constant pose while the carriers move because it is obviously the most challenging case compared to the one in which also the load moves.
The work is organized as follows. First, in Sec. II, we introduce the notation and we state the problem formally. Hence, we answer the above-described research question in Sec. III, in which we analyze the cases of one, two, and three flying carriers. We show that 3 is the minimum number of flying carriers for which non-stop flights are possible and we derive the conditions for that to hold. In Sec. IV we show numerical examples in support of the theoretical results. Eventually, conclusions are drawn in Sec. V with an outline of future work.
II Problem Statement
II-A Mathematical background
Consider ${\mathcal{F}_{W}=\{O_{W},\bm{x}_{W},\bm{y}_{W},\bm{z}_{W}\},}$ being a world-fixed frame where $O_{W}$ is the origin, and $\bm{x}_{W},\bm{y}_{W},\bm{z}_{W}$ are the x-, y-, and z- axes, respectively. Similarly, consider a frame attached to the Center of Mass (CoM) of a rigid body, which represents our manipulated object, and call it ${\mathcal{F}_{B}=\{O_{B},\bm{x}_{B},\bm{y}_{B},\bm{z}_{B}\}.}$ The position of $O_{B}$ in $\mathcal{F}_{W}$ is denoted with ${\bm{p}_{L}\in\mathbb{R}^{3}}$ and $\bm{R}_{L}\in SO(3)$ expresses the attitude of $\mathcal{F}_{B}$ w.r.t $\mathcal{F}_{W}$. The body angular velocity expressed in the body frame is indicated as ${}^{B}\bm{\omega}_{L}$.
The rigid body is manipulated through $n$ cables, each one of length $L_{i}>0$ and attached to point $B_{i}$ on the object, with ${}^{B}\bm{b}_{i}\in\mathbb{R}^{3}$ representing the constant position of $B_{i}$ expressed in $\mathcal{F}_{B}$ [^1] The left superscript expresses the reference frame, and if omitted it is considered equal to $W$ unless differently specified.. Let $\bm{q}_{i}\in S^{2}$ be the configuration of the $i^{\rm{th}}$ cable in $\mathcal{F}_{W}$, $T_{i}$ its tension, and
$\bm{f}_{i}=T_{i}\bm{q}_{i}$ (1)
be the coordinates in $\mathcal{F}_{W}$ of the force that the $i^{\rm{th}}$ cable exerts on the load. The cables’ mass and inertia are considered negligible compared to the robots’ and object’s ones, and cables are considered in tension under the effect of the weight [21, 14] (and so $T_{i}\neq 0)$. Each cable is attached, at the other end, to a point on an aerial robot (flying carrier), whose position in $\mathcal{F}_{W}$ is denoted with $\bm{p}_{Ri}\in\mathbb{R}^{3}$.
The equations describing the dynamics of the transported object are
${m_{L}}\ddot{\bm{p}}_{L}$ $=-{m_{L}}g\bm{e}_{3}+\sum_{i=1}^{n}{\bm{f}_{i}}$ (2)
$\bm{J}_{L}{{}^{B}\dot{\bm{\omega}}}_{L}$ $=-\bm{J}_{L}{{}^{B}\bm{\omega}}_{L}\times{{}^{B}\bm{\omega}}_{L}+\sum_{i=1}^{n}{S(^{B}\bm{b}_{i})\bm{R}_{L}\bm{f}_{i}}$ (3)
$\dot{\bm{R}}_{L}$ $=S(\bm{\omega}_{L})\bm{R}_{L}$ (4)
where ${m_{L}},\bm{J}_{L}$ are the mass and the rotational inertia of the rigid body load and $S(\star)$ indicates the skew operator that implements the cross product between two vectors.
The position and velocity of the $i^{\rm{th}}$ flying carrier in $\mathcal{F}_{W}$ are, from the system’s kinematics, equal to
$\bm{p}_{Ri}$ $=\bm{p}_{L}+\bm{R}_{L}{{}^{B}\bm{b}}_{i}+\bm{q}_{i}L_{i}$ (5)
$\dot{\bm{p}}_{Ri}$ $=\dot{\bm{p}}_{L}+\dot{\bm{R}}_{L}{{}^{B}\bm{b}}_{i}+\dot{\bm{q}}_{i}L_{i}.$ (6)
To compactly rewrite the load dynamics we use the matrix $\bm{G}\in\mathbb{R}^{6\times 3n}$ that maps the cable forces to the wrench applied at the object’s center of mass, referred to as grasp matrix in the literature [22, 23]
$\bm{G}=\begin{bmatrix}\bm{I}_{3}&\bm{I}_{3}&...&\ \bm{I}_{3}\\ S(^{B}\bm{b}_{1})\bm{R}_{L}&S(^{B}\bm{b}_{2})\bm{R}_{L}&...&S(^{B}\bm{b}_{n})\bm{R}_{L},\end{bmatrix}.$ (7)
The load dynamics (2)-(3) can be compactly rewritten as
$\bm{W}=\bm{G}\bm{f}$ (8)
where $\bm{f}=[\bm{f}_{1}^{\top}\ \bm{f}_{2}^{\top}\ ...\ \bm{f}_{n}^{\top}]^{\top}$ stacks all forces that the cables apply to the object and ${\bm{W}=\begin{bmatrix}{m_{L}}(\ddot{\bm{p}}_{L}+g\bm{e}_{3})\\ \bm{J}_{L}{{}^{B}\dot{\bm{\omega}}}_{L}+\bm{J}_{L}{{}^{B}\bm{\omega}}_{L}\times{{}^{B}\bm{\omega}}_{L}\end{bmatrix}}$.
Inverting (8), see [24], we obtain an expression for all the cable forces $\bm{f}$ that are compatible with a certain $\bm{W}$
$\bm{f}=\bm{G}^{\dagger}\bm{W}+\bm{N}\bm{\lambda},$ (9)
with $\star^{\dagger}$ indicating the Moore-Penrose pseudo-inverse, $\bm{N}\in\mathbb{R}^{3n\times m}$ being the nullspace projector of the grasp matrix $\bm{G}$ (with $m$ the dimension of the nullpsace of $\bm{G}$), and $\bm{\lambda}\in\mathbb{R}^{m}$ a vector of $m$ free parameters $\lambda_{1},\ldots,\lambda_{m}$ referred to as the internal forces [22].
For each load forced equilibrium or trajectory defined by the load state and the corresponding forcing input $\bm{W}$ we can obtain from (9) all the cable forces that are compatible with such load equilibrium/trajectory. These forces are parameterized by $\bm{\lambda}$. In turn, to each cable force corresponds a carrier position obtained by (5) substituting $\bm{q}_{i}=\bm{f}_{i}/T_{i}$. So we can conclude that for any given load equilibrium/trajectory there are $\infty^{m}$ compatible multi-UAV coordinated trajectories which are parameterized by the internal forces $\bm{\lambda}$ and are obtained by plugging (9) in (5).
II-B Maintaining Load Static Equilibrium with non-stop flights
Internal forces have been exploited in the literature of cooperative aerial manipulation of suspended objects to, e.g., optimize the tension distribution in the carriers’ cables [19], perform obstacle avoidance [25], or stabilize the object’s pose in a force-based manipulation scenario [14]. However, all the aforementioned works consider multirotor UAVs that can stop (i.e., set $\dot{\bm{p}}_{Ri}=\bm{0}$) during the manipulation task execution.
Departing from the previous literature, in this work we explore whether the internal forces can be exploited to allow the flying carriers to keep loitering with $\|\dot{\bm{p}}_{Ri}\|\geq\underline{v}>0$ while the load pose is kept to a static pose.
Consider a static equilibrium of the suspended load characterized by position $\bar{\bm{p}}_{L}$ and orientation $\bar{\bm{R}}_{L}$. When the load is at a static equilibrium, the only way the cable forces can still change is thanks to a change of the internal forces $\bm{\lambda}$ since they are the only time-varying parameters in (9) in that case. For the purpose of analyzing the effect of changing internal forces we write them as the state of the following dynamical system:
$\begin{cases}\dot{\bm{\lambda}}=\bm{u}_{\lambda}\\ \bm{\lambda}(0)=\bm{\lambda}_{0}\end{cases}$
where $\bm{u}_{\lambda}$ acts as an input that we can assign and $\bm{\lambda}_{0}$ is their initial condition.
Problem 1 (pose regulation with non-stop flights).
Assume that $\bm{p}_{L}(0)=\bar{\bm{p}}_{L}$ and $\bm{R}_{L}(0)=\bar{\bm{R}}_{L}$. We ask if there exists a $n$ and a trajectory $\bm{\lambda}(t)$ such that $\forall t\geq 0,$
$\begin{cases}\bm{p}_{L}(t)=\bar{\bm{p}}_{L},\\ \bm{R}_{L}(t)=\bar{\bm{R}}_{L},\\ \|\dot{\bm{p}}_{Ri}(t)\|\geq\underline{v}>0,\quad\forall i=1,\ldots,n\\ -\infty<\underline{\lambda}\leq\lambda_{i}(t)\leq\overline{\lambda}<\infty,\quad\forall i=1,\ldots,n\end{cases}$ (10)
The first two equations in (10) impose static equilibrium of the load. The third equation, where $\underline{v}$ is a constant, is a lower bound to the norms of the flying carriers’ velocities to impose that they do not stop while maintaining the constantly desired pose of the object. Finally, the last expression in (10) imposes that each component of $\bm{\lambda}$ is bounded by certain upper and lower bounds, $\overline{\lambda},\underline{\lambda}$. This last constraint has clear practical reasons as internal forces of too large magnitude in (9) generate cable forces that will eventually break the system.
When (10) holds, (9) can be rewritten component-wise as
$\bm{f}_{i}(t)=\bm{f}_{0i}+\tilde{\bm{f}}_{i}(t),$ (11)
where $\bm{f}_{0i}$ is the component of each force that comes from the first term in the left-hand side (LHS) of (9), namely $\bm{G}^{\dagger}\bm{W}$; it balances the external wrench and is constant when the load is static. Instead, $\bm{\tilde{f}}_{i}$ is the variable component of the cable force generated by the internal forces, which comes from $\bm{N}\bm{\lambda}(t)$ in (9).
Taking the time derivative of (9) we obtain:
$\dot{\bm{f}}={\bm{N}}\dot{\bm{\lambda}}.$ (12)
Equation (12) gives us a relationship between the internal force variation and the total cable force variation; to give an intuition, it gives all the possible variations of the cable forces, and hence of the carriers’ velocities, that do not perturb the object pose equilibrium.
By differentiating (1) w.r.t. time, the cable force variation (12) is also
$\dot{\bm{f}}_{i}=\dot{T}_{i}\bm{q}_{i}+T_{i}\dot{\bm{q}}_{i},$ (13)
i.e., the force variation is composed of two orthogonal components: one is directed as $\bm{q}_{i}$ (along the cable), and one is $\dot{\bm{q}}_{i}$ (orthogonal to the cable).
III Non-stop Flight Analysis
In this section, we will reply to the question: ‘are non-stop carrier UAV flights compatible with keeping cable-suspended objects at a static pose?’, which is formally stated in Problem 1.
First, we introduce here a few tools to reply to the above-formulated research question. We start by formalizing an intuitive condition.
Proposition 1.
When the load is at a static equilibrium and the cables are in tension,
$\exists\underline{v}>0:\|\dot{\bm{p}}_{Ri}\|\geq\underline{v}\;\forall t\Longrightarrow\exists s>0:\|\dot{\bm{f}}_{i}\|\geq s\;\forall t$
Proof.
By computing (6) at the load’s static equilibrium, we have
$\dot{\bm{p}}_{Ri}=\dot{\bm{q}}_{i}L_{i}.$ (14)
Hence, by using $\|\dot{\bm{p}}_{Ri}\|\geq\underline{v}$ inside (14) and given that the cable length $L_{i}$ is greater than zero, we obtain
$\|\dot{\bm{q}}_{i}\|\geq\frac{\underline{v}}{L_{i}}.$ (15)
Consider now that, from (13), the cable forces vary along two orthogonal directions. That, plus the fact that $\bm{q}_{i}$ is a unit vector by definition, leads to
$\|\dot{\bm{f}}_{i}\|=|\dot{T}_{i}|+|T_{i}|\|\dot{\bm{q}}_{i}\|$ (16)
Now, starting from (13) and using (15), (16), $L_{i}>0$ and $|T_{i}|\neq 0$ by hypothesis of tensioned cables, we write:
$\|\dot{\bm{f}}_{i}\|=|\dot{T}_{i}|+|T_{i}|\|\dot{\bm{q}}_{i}\|\geq|\dot{T}_{i}|+|T_{i}|\frac{\underline{v}}{L_{i}}\geq|T_{i}|\frac{\underline{v}}{L_{i}}:=s>0.$
Remark 1.
Proposition 1, when looked in reverse, provides a necessary condition for (10) to hold. Namely, in order to have all the carriers perform non-stop flights, all the norms of the cable forces derivatives must be lower bounded (i.e., if $\forall s>0,\ \|\dot{\bm{f}}_{i}(t)\|\ngeq s\Rightarrow\forall\underline{v}>0,\ \|\dot{\bm{p}}_{Ri}\|\ngeq\underline{v}$).
III-A 1-carrier case
First, we consider the case in which the is only one flying carrier, i.e., it is $n=1$. The sole equilibrium points of the system are in this case the ones in which the load’s center of mass lies below the vertical cable that suspends it [^2] Despite this number of carrier(s) does not allow regulating the full pose of a suspended object, we are here interested in understanding if the carrier(s) can loiter while the object is at a static equilibrium..
The grasp matrix in (7) is
$\bm{G}=\begin{bmatrix}\bm{I}_{3}\\ S(^{B}\bm{b}_{1})\bm{R}_{L}\end{bmatrix}.$
Hence, as the grasp matrix $\bm{G}\in\mathbb{R}^{6\times 3}$ has number of columns smaller than the number of rows, it has a trivial nullspace and $\bm{N}=\bm{0}$.
Fact 1.
For $n=1$, $\nexists t$ such that conditions (10) hold.
Proof.
$\bm{N}=\bm{0}$ implies that (12) becomes $\dot{\bm{f}}\equiv\bm{0}$. According to (13), the cable force varies along two orthogonal components. Hence, the only way in which $\dot{\bm{f}}_{i}=\bm{0}$ is that both those orthogonal components are equal to zero. In other words, both $\dot{T}_{i}=0$ and $\dot{\bm{q}}_{i}=\bm{0}$ (because $T_{i}\neq 0$ by assumption). When the load is at equilibrium, $\dot{\bm{p}}_{L}=\bm{0}$ and $\dot{\bm{R}}_{L}=\bm{0}$. By applying all the above conditions to (6), one obtains $\dot{\bm{p}}_{Ri}=\bm{0}$. ∎
In conclusion, Fact 1 tells that for $n=1$ non-stop flights for maintaining the static equilibrium of a suspended object are not possible.
III-B 2-carrier case
[…]
$\bm{G}=\begin{bmatrix}\bm{I}_{3}&\bm{I}_{3}&\\ S(^{B}\bm{b}_{1})\bm{R}_{L}&S(^{B}\bm{b}_{2})\bm{R}_{L}&\end{bmatrix}.$
$\bm{G}\in\mathbb{R}^{6\times 6}$ has a non-trivial nullspace of dimension $m=1$ and a basis of if is built from the vector along the direction through the points $B_{1}$ and $B_{2}$ [14]. We indicate the vector that constitutes such basis with $\bm{n}_{1}$, and $\lambda_{1}$ the corresponding scalar internal force. In particular, $\bm{N}={\bm{n}_{1}=\begin{bmatrix}\bm{b}_{1,2}^{\top}&\bm{b}_{2,1}^{\top}\end{bmatrix}^{\top}},$ where we indicate in general the unit vector in the direction $B_{i}B_{j}$ as ${\bm{b}_{i,j}=\frac{\bm{b}_{i}-\bm{b}_{j}}{|\bm{b}_{i}-\bm{b}_{j}|}}$.
Fact 2.
For $n=2$, $\nexists\bm{\lambda}(t)$ such that $\forall t$ (10) holds.
Proof.
At the load equilibrium, $\bm{n}_{1}$ is constant by construction. Equation (12) leads to
$\dot{\bm{f}}_{i}=\dot{\lambda}_{1}{\bm{b}_{i,j}}.$ (17)
Assume that (10) holds. Then, for Proposition 1, $\exists s>0:\|\dot{\bm{f}}_{i}\|\geq s$. For the considered case of 2 carriers, $\forall t>0,$ $\exists s>0:\|\dot{\bm{f}}_{i}\|\geq s\Longleftrightarrow|\dot{\lambda}_{1}|\geq s.$ This follows from (17) considering that $\bm{b}_{i,j}$ has unit norm by construction. However, $|\dot{\lambda}_{1}|\geq s$ implies that $|\lambda_{1}|$ is a strictly increasing function of time and therefore is unbounded for $t\to\infty$. This contradicts the boundedness of $\lambda_{1}$ imposed in (10) and therefore (10) cannot hold as two of its conditions are in conflict. ∎
In conclusion, Fact 2 tells us that non-stop UAV flights for the pose regulation of a cable-suspended object manipulated by 2 carriers are not possible.
III-C 3-carrier case
In this section, we consider the case in which $n=3$ carriers are employed. Let us restrict the analysis to the non-degenerate cases in which the cable attachment points on the objects, $B_{i}$, are not all aligned to each other. In that degenerate case, the problem is similar to the 2-carrier scenario. In our study, we are going to assume that the minimum angle between two vectors $\bm{b}_{i,j}$ is $|\theta_{i,j}|\leq\bar{\theta}_{i,j}>0$.
For $n=3$, $\bm{\lambda}$ is a 3-dimensional vector of components $\lambda_{i}$, for $i\in\{1,2,3\}$, and a choice for $\bm{N}$ is expressed starting from the directions connecting the three points $B_{i}$ in pairs. In particular, it holds that [24, 22]:
$\bm{N}=\begin{bmatrix}\bm{b}_{1,2}&\bm{0}&\bm{b}_{1,3}\\ -\bm{b}_{1,2}&\bm{b}_{2,3}&\bm{0}\\ \bm{0}&-\bm{b}_{2,3}&-\bm{b}_{1,3}\end{bmatrix}.$ (18)
Equation (18) tells us that the components of each cable force generated by the internal forces are a couple of equal and opposite forces acting along the line connecting two cable attachment points.
Figure 2: Geometric representation of the cable force, its components, and its derivative. The grey vectors $\dot{\bm{f}}_{i}$ live in plane $P_{i}$, while its components $T_{i}\dot{\bm{q}}_{i},\dot{T}_{i}\bm{q}_{i}\notin P_{i}$ (in general). $\tilde{\bm{f}}_{i}$ is contained on the subspace $\tilde{P}_{i}$, which is centered at the end-point of $\bm{f}_{0i}$ and contains also the end-point of $\bm{f}_{i}$.
Figure 3: Representation of the plane including $\dot{\bm{f}}_{i}$ and its components along the directions parallel and perpendicular to the corresponding cable force. The minimum and maximum values of the angle between $\dot{\bm{f}}_{i}$ and $\bm{f}_{i}$ are represented with dotted incident lines, and the circle of radius $s$ is the lower bound on the norm of $\dot{\bm{f}}_{i}.$ As a consequence, the norm of the component of $\dot{\bm{f}}_{i}$ on the direction orthogonal to the cable force, namely along $\dot{\bm{q}}_{i}$, is lower bounded by a positive quantity, $s^{\prime\prime}$.
Figure 4: Two carriers cannot maintain the pose of the object constant while performing non-stop flights. On the left, is a screenshot of an animation: the object is represented as a grey line, the cables are in black, and colored curves are the carrier’s paths. Carriers’ velocities and $\dot{\bm{\lambda}}$ are reported in the other two plots.
Proposition 2.
For $n=3$, if the load’s attitude at the static equilibrium is such that $\forall i$ $\bm{f}_{0i}\notin{\rm{span}}\{\bm{b}_{i,j},\bm{b}_{i,k}\},$ with $i,j,k\in\{1,2,3\},$ $i\neq j\neq k$, then for $s,\underline{v}>0$, $\|{\dot{\bm{f}}_{i}\|\geq s\Rightarrow\|\dot{\bm{p}}_{Ri}\|\geq\underline{v}.}$ $\forall i=1,\ldots,n$
Proof.
First of all, from (6) at the equilibrium, ${\|\dot{\bm{p}}_{Ri}\|=L_{i}\|\dot{\bm{q}}_{i}\|}$; hence, to prove the thesis of the proposition it is enough to show that $\|\dot{\bm{q}}_{i}\|\geq s^{\prime}$ with $s^{\prime}>0$. For the sake of a streamlined presentation we decided to avoid presenting all the mathematical details and to make use of explanatory figures that help the intuition of the reader being convinced about the validity of the proof.
Please refer to to Fig. 2 to better understand the quantities defined in the following. Referring to (11) we know that ${\bm{f}_{i}=\bm{f}_{0i}+\tilde{\bm{f}}_{i}}$, where $\bm{f}_{0i}$ is obtained by the corresponding entries of $\bm{G}^{\dagger}\bm{W}$ and it is nonzero and constant in time, and $\tilde{\bm{f}}_{i}$ is obtained by the corresponding entries of $\bm{N}\bm{\lambda}$ and it is time-varying and belongs to the plane $P_{i}$ spanned by the vectors $\bm{b}_{i,j}$ and $\bm{b}_{i,k}$ with $j,k\neq i$ and $j\neq k$, according to (18). From the hypotheses of the proposition, we have that $\bm{f}_{0i}$ does not belong entirely to $P_{i}$ (apart from its endpoint). Since $\bm{f}_{0i}$ is constant and nonzero, also $\bm{f}_{i}$ does not belong entirely to $P_{i}$ (apart from its endpoint). On the other hand, from (12), we have that $\dot{\bm{f}}_{i}$ belongs entirely to $P_{i}$. As a consequence, thanks also to the hypothesis that $\dot{\bm{f}}_{i}\neq\bm{0}$, we obtain that $\dot{\bm{f}}_{i}$ and $\bm{f}_{i}$ are never parallel to each other (two parallel vectors must belong to the same plane). In other words, the angle between each cable force and its derivative is $0<\angle\bm{f}_{i},\dot{\bm{f}}_{i}<\pi.$
Because of the boundedness of each component of $\bm{\lambda}$, the endpoint of $\bm{f}_{i}$ is constrained inside a bounded set denoted with $\tilde{P}_{i}\in P_{i}$, the dark grey region in Figure 2. $\tilde{P}_{i}$ is centered around the end-point of $\bm{f}_{0i}$ and contains the vector $\tilde{\bm{f}}_{i}.$ For all the $\bm{f}_{i}$, we have a minimum and maximum of the angle $\angle\bm{f}_{i},\dot{\bm{f}}_{i}$ when $\dot{\bm{f}}_{i}$ varies in $\tilde{P}$, which we denote as $\alpha(\bm{f}_{i})$ and $\beta(\bm{f}_{i})$ respectively. Furthermore, we denote with $\underline{\alpha}$ and $\overline{\beta}$ the minimum and maximum of of $\alpha(\bm{f}_{i})$ and $\beta(\bm{f}_{i})$ when the end point of $\bm{f}_{i}$ varies in $\tilde{P}_{i}$, respectively. Since those functions are continuous and the set $\tilde{P}_{i}$ is bounded such minimum and maximum exist and therefore, for what has been previously said, they are such that $0<\underline{\alpha}\leq\overline{\beta}<\pi.$
Now let’s consider the two orthogonal components of $\dot{\bm{f}}_{i}$ defined in (13). These vectors, together with $\dot{\bm{f}}_{i}$ and $\bm{f}_{i}$ belong all to a plane that is represented in Figure 3. In that figure, the bounds of the angles between $\dot{\bm{f}}_{i}$ and the direction of the force $\bm{f}_{i}$, given by $\bm{q}_{i}$, are highlighted by incident lines properly sloped. Moreover, the lower bound of the norm of the force derivative, $s>0$, is represented by a dotted circle; the endpoint of $\dot{\bm{f}}_{i}$ is constrained to be outside of that circle. Hence, its component on the direction orthogonal to the force, namely the one parallel to $\dot{\bm{q}}_{i}$, must lie in the green regions in Figure 3. This gives a geometrical proof that $\|\dot{\bm{q}}_{i}\|T_{i}\geq s^{\prime\prime}>0$ and so $\|\dot{\bm{q}}_{i}\|\geq\frac{s^{\prime\prime}}{T_{i}}:=s^{\prime}>0.$
We are now ready to state the main result.
Fact 3.
[…]
$\lambda_{i}=\lambda_{i}(0)+A\cos{(\psi t+\Phi_{i})}$ (19)
with constant amplitude $0<A<\infty$, initial condition $\underline{\lambda}\leq\lambda_{i}(0)\leq\overline{\lambda}$, constant phases $\Phi_{i}$ different from each other such that $|\Phi_{i}-\Phi_{j}|\geq\phi>0$ and such that $0<\underline{\Phi_{i}}\leq\Phi_{i}\leq\bar{\Phi}<\pi,$ and same positive $\psi$.
Proof.
Let us consider without loss of generality the force applied by cable 1. From equation (12),
$\dot{\bm{f}}_{1}=\dot{\lambda}_{1}\bm{b}_{1,2}+\dot{\lambda}_{3}\bm{b}_{1,3}.$ (20)
When ${\lambda}_{1}$ and ${\lambda}_{3}$ are chosen as in (19), they are bounded because the initial conditions and amplitudes $A$ are bounded. Especially, $\forall t,$ $|\lambda_{i}(t)|\leq|\lambda_{i}(0)|+|A|$. Moreover, their derivatives are $\dot{\lambda}_{i}=-A\psi\sin{(\psi t+\Phi_{i})}$. Because we assumed $\bm{b}_{i,j}$ are not parallel to each other, for $\|\dot{\bm{f}}_{i}\|$ to be lower bounded it is sufficient that at each time at least one of its components along the non-parallel vectors $\bm{b}_{i,j}$ is lower bounded. For the hypothesis on $\Phi_{1}$ and $\Phi_{2}$, the two terms in the RHS of (20) are such that $\forall t,$ $|\dot{\lambda}_{1}|\geq\bar{\dot{\lambda}}_{1}>0$ or ${|\dot{\lambda}|_{2}\geq\bar{\dot{\lambda}}_{2}>0}$. In the worst-case scenarios, one of the components in the RHS of (20) is zero, and $\|\dot{\bm{f}}_{i}\|=\bar{\dot{\lambda}}_{i}>0.$ Using Proposition 2, we conclude that the intensities of the carriers’ velocities are lower bounded, too. ∎
Remark: Fact 3 positively answer Problem 1 telling us that pose regulation of a suspended object using non-stop flying carriers is possible for $n=3$.
(a) Three flying carriers execute non-stop flights while the suspended object pose is kept at a constant pose.
(b) One of the carriers stops during its motion because two components $\dot{\lambda}_{i}$ that generate its cable force are equal to each other in time.
(c) One of the carriers stops during its motion when the two components $\dot{\lambda}_{i}$ that generate its cable force vanish simultaneously.
Figure 5: Three different simulations. On the leftmost column, non-stop flights are shown; in the other two columns, cases in which the carriers stop. On the first row, screenshots from tasks in which the 3 flying carriers maintain the suspended object in a constant pose. The area of the object among the cable attachment points is highlighted with dotted lines, the cables are black lines, and the carriers at their endings are black dots; the path of each carrier is the colored line in blue for carrier 1, red for carrier 2, and yellow for carrier 3; the cross is the object CoM. The values of $\|\dot{\bm{p}}_{Ri}\|$ (mid row) and $\dot{{\lambda}_{i}}$ (bottom row) are reported in the corresponding columns.
Even though we showed that it is possible for 3 flying carriers to maintain a constant pose of suspended objects without stopping, it is also interesting to analyze special choices of $\dot{\bm{\lambda}}$, if any, that do not generate non-stop flights even for 3 carriers.
3 carriers—degenerate case 1
If two components of $\dot{\bm{\lambda}}$, let us consider without loss of generality $\dot{\lambda}_{1}$ and $\dot{\lambda}_{2}$, are such that $\dot{\lambda}_{2}=a\dot{\lambda}_{1}$, $a\in\mathbb{R}$, then the corresponding force variation, in this case as in (20), becomes
$\dot{\bm{f}}_{1}=(a+1)\dot{\lambda}_{1}(\bm{b}_{1,2}+\bm{b}_{1,3}),$ (21)
which implies that $\dot{\bm{f}}_{1}$ lies along a constant direction over time, similar to what happens for the two-carrier case. We recall that $\bm{b}_{i,j}$ are unit vectors and are not parallel to each other by hypothesis. Hence, $|\bm{b}_{1,2}+\bm{b}_{1,3}|>0$ and this case can be conducted to the two-carrier case. So, for the same reasoning of the proof of Fact 2, the velocity of carrier 1 is not lower bounded for all time, and so (10) does not hold.
3 carriers—degenerate case 2
In any case in which $\exists t:\dot{\lambda}_{i}(t)=\dot{\lambda}_{j}(t)=0$ for some $i\neq j\in\{1,2,3\}$ then the corresponding force is zero at that time instant. For the same reasoning done in the proof of Fact 1, the corresponding carrier stops. Differently from the situation described above, the cable force variation generally is not on a straight line.
IV Numerical Results
In this section, we propose simulation results that support the previously presented results.
The simulations have been carried out in Matlab-Simulink. The dynamics of each carrier is a double integrator controlled through a PD feedback law to follow a desired trajectory. Indeed, the scope of this work is to assess the feasibility of non-stop flights for the considered application. Implementation of such flights considering more realistic non-stop UAV dynamics is left for future work.
The carriers’ mass is 0.1 kg, and the derivative and proportional gain matrices are diagonal with values on the diagonals equal to 1.5 Ns/m and 1000 N/m, respectively.
[…]
Each internal force is chosen as in (19), with parameters given in the following. The corresponding cable force variations are computed as in (12) and the carriers’ trajectories from (5) and (6).
[…]
Figure 5b and 5c show an example of what was referred to as degenerate case 1 and 2, respectively, in which 3 carriers maintain the object’s pose but one of them stops a few times along the path. In the simulation of case 1, $\forall i,\ \lambda_{i}(0)=2$ N $A=1$ N $\psi=1$ rad/s, $\Phi_{1}=\Phi_{3}=0$ and $\Phi_{2}=0.7$ rad. In the simulation of case 2, instead, $\forall i,\ \lambda_{i}(0)=2$ N, $A=1.2$ N, $\psi_{1}=\psi_{3}=1$ rad/s, $\psi_{2}=0.5$ rad/s, $\Phi_{1}=\Phi_{2}=0$, and $\Phi_{3}=0.7$ rad. The carriers’ velocities are reported in Figures 5e and 5f for the two cases, respectively. As it can be seen in Figures 5h and 5i, for some time instants $\dot{\lambda}_{1}=\dot{\lambda}_{j}=0$. At the same time instants, the velocity of one carrier goes to zero, too, as predicted by the theory.
V Conclusions and Future Work
This work studied the compatibility of non-stop carrier flights while keeping constant the pose of cable-suspended objects. After showing that non-stop flights are not feasible when only one or two flying carriers are used, we showed that three carriers can follow non-stop flights while maintaining the object’s pose constant. We also highlighted degenerate cases in which non-stop flight is not achieved even for three carriers. Eventually, numerical examples are shown to support the theoretical analysis.
In the future, we will extend the study to a generic number of carriers $n>0$. Also, we will design an optimal planning algorithm that generates smooth non-stop trajectories for UAVs, explicitly taking into account the system’s constraints, such as the minimum forward speed of the robots, their maximum angular velocity, collision avoidance, etc. Keeping the forward speeds of the robots as constant as possible could also be considered in view of energy efficiency. More realistic simulations will be developed to assess the practical applicability of the manipulation method, e.g., considering hybrid VTOL vehicles.